OpenStudy (anonymous):

Verify the identity. tan(x+(pi/2)) = -cot x

OpenStudy (zzr0ck3r):

tan(x+pi/2) = sin(x+pi/2)/cos(x+pi/2) = cos(x)/-sin(x) = -cot(x)

OpenStudy (zzr0ck3r):

understand?

OpenStudy (zzr0ck3r):

sin(x+pi/2) = cos(x) and cos(x+pi/2) = -sin(x)

OpenStudy (anonymous):

Somewhat. I understand the identities well, just a bit unsure as to how tan(x+pi/2) can =sin(x+pi/2)/cos(x+pi/2). I know that tan=sin/cos but why did all the ()carry?

OpenStudy (zzr0ck3r):

tan(anything) = sin(anything)/cos(anything) so tan(2) = sin(2)/cos(2) tan(2x) = sin(2x)/cos(2x) tan(puppy dog) = sin(puppy dog)/cos(puppy dog) so tan(x+pi/2) = sin(x+pi/2)/cos(x+pi/2)

OpenStudy (anonymous):

ok, I thought just the theta or x in this case would follow. Thanks so much

OpenStudy (zzr0ck3r):

nope: everything inside the () :)